OpenStudy (pigsruleyeah):

1 over 3b = 4 over 5. Which of the following equals b in this equation? 22 over 5 11 over 8 2 over 5 1 over 4

RhondaSommer (rhondasommer):

\[\frac{1}{3b}=\frac{4}{5}\]

RhondaSommer (rhondasommer):

cross multiply :D \[3b*4=1*5\]can you simplify that?

OpenStudy (peytonrychtik):

c